OpenStudy (anonymous):

Complete the square to find the vertex of each quadratic function. State the vertex. y=5x^2+1/5x−10

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

y=5(x^2+1/25 x)-10 y=5{x^2+1/25 x+(1/50)^2-(1/50)^2}-10 y=5(x+1/50)^2-5*1/2500 -10 y==5(x+1/50)^2-5001/500

OpenStudy (anonymous):

oh thank you so much !

OpenStudy (anonymous):

yw